# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ Abstract
A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures. This paper introduces a decentralized decoding architecture that fundamentally breaks from the conventional LSFD, by allowing each AP calculates interference-suppressing local weights independently and applies them to its data estimates before transmission. Furthermore, two generalized local zero-forcing (ZF) framework, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each access point (AP) adaptively and independently determines its combining strategy through a local sum spectral efficiency optimization that classifies user equipments (UEs) as strong or weak using only local information, eliminating the fixed thresholds used in PFZF and PWPFZF. To further enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot. The corresponding closed-form spectral efficiency expressions are derived. Numerical results show that the proposed generalized schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overhead and computation costs with minimal performance penalty compared to them.
@@ I Introduction
Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee consistent service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. Unlike conventional cellular architectures, such distributed systems enhance coverage and reliability through cooperative processing and coherent transmission/reception [3]. To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [4]. A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
@@
Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs. Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments. To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE. Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels. These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios. These centralized and partial centralized combining schemes follow the the fully centralized architecture as shown in Fig. 1, which has very high computational cost and fronthaul overhead, as for all UEs, channel need to be estimated at the CPU with the cooperation of all or a subset of APs.
@@
Several distributed combining approaches have been proposed in the literature. Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference. To improve upon MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance. However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived. As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult to obtain. In [1, 3], local combining schemes follow an architecture of simple decoding as shown in Fig. 1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality.
(a) FullyCentralized(b) Local Combining and LSFD(c) Local Combining and Simple decoding (d) Proposed Decentralized decoding APs CPU Received SignalsChannel Estimates Global Data Estimates $\hat{x}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$Local Channel Estimates Calculates LSFD Weights $\mathbf{a}_{t}$ Final Estimates $\hat{x}_{t}=\mathbf{a}_{t}^{H}\hat{\mathbf{x}}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$ Final Estimates $\hat{x}_{t}=\sum_{m}\hat{x}_{mt}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ Local Weighting $a_{mt}\cdot\hat{x}_{mt}$ CPU Pre-weighted DataEstimates $a_{mt}\hat{x}_{mt}$ Final Data Estimates $\hat{x}_{t}=\sum_{m}a_{mt}\hat{x}_{mt}$
Figure 1: Comparison of uplink processing architectures in D-mMIMO: (a) Fully centralized processing is computationally prohibitive. (b) Local combining and LSFD requires sharing soft estimates and global CSI to compute optimal weights at the CPU. (c) Local combining and Non-coherent decoding is simple but sub-optimal. (d) The proposed decentralized decoding architecture: each AP performs local combining and weighting independently, reducing the CPU to a simple aggregator and eliminating fronthaul overhead for coordination.
@@
The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form spectral efficiency (SE) expressions. These techniques construct combining vectors locally at each AP using channel statistics to suppress interference within designated user groups. In traditional FZF, PFZF, and PWPFZF, each AP generates a single combining vector per orthogonal pilot sequence and assigns scaled versions of this vector to individual UEs based on their estimated channel variances [13]. The fundamental limitations of PFZF and PWPFZF stem from two factors. First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources. Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
@@
The large-scale fading decoding (LSFD) as shown in Fig. 1, is employed at the central processing unit (CPU) to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14]. In the optimal LSFD (o-LSFD) approach [13], each AP must transmit both channel estimates and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16]. To alleviate this burden, partial LSFD (p-LSFD) is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process. However, p-LSFD still requires matrix inversions of the same dimension as o-LSFD and the transmission of channel estimates to the CPU. As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of LSFD-based architectures in ultra-dense D-mMIMO deployments.
@@ I-A Contributions
To overcome the above mentioned issues, this work introduces a novel decentralized decoding framework for D-mMIMO that fundamentally enhances scalability, adaptability, and analytical tractability. The principal contributions are summarized as follows:
@@
  • Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig. 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead. Central to this architecture is the design of interference-suppressing local weights computed independently by each AP using only its local channel information. These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator. This eliminates the need for any inter-AP coordination or channel information sharing with the CPU, reducing it to a low-complexity aggregator. This transition from a coordinated to a truly distributed architecture is the key enabler for the practical, large-scale deployment of cell-free networks.
@@
  • Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Unlike conventional PFZF and PWPFZF, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy. Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR. Thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
@@
  • Novel Distributed Local Optimization for Pilot Partitioning: We also propose novel distributed optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum spectral efficiency metric. This pilot-level optimization is solved using projected gradient ascent optimization and therefore maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
@@
  • Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable spectral efficiency of the proposed system. This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
@@
  • Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good performance with reduced computation cost with state-of-art baselines. The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul and computational costs in comparison to LSFD framewrok and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
@@
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions. Section IV provides numerical simulations to evaluate performance. Finally, Section V concludes the paper and outlines potential directions for future work.
@@
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The cardinality of a set $\mathcal{S}$ is denoted by $|\mathcal{S}|$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
@@ II System Model
The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas, ensuring that $T\ll MA$. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources. Every AP is linked to the CPU by fronthaul link, so that the network can operate together and process signals. Using uplink pilot transmissions, the system estimates the AP-UE channels under the time-division duplexing (TDD) regime. The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for uplink pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The comprehensive channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ between the AP $m$ and the UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$. We suppose that $\textbf{h}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\beta_{mt}\textbf{I}_{A})$.
@@ II-A Channel Estimation
A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of the UE $t$. At the $m$-th AP, the received signal $\textbf{y}^{\text{pilot}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is provided by:
$\textbf{y}^{pilot}_{m}=\sum_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized signal-to-noise power ratio for the UE $t$ during pilot transmission. The minimum mean square error (MMSE) estimate $\hat{\mathbf{g}}_{mt}\in\mathbb{C}^{A\times 1}$ of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=c_{mt}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
where
${c}_{mt}=\frac{\sqrt{p^{p}_{t}L_{p}}\beta_{mt}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}.$
The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}={\mathbf{g}}_{mt}-\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,(\beta_{mt}-\gamma_{mt})\textbf{I}_{A})$, where
$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=\left(\frac{p^{p}_{t}L_{p}\beta_{mt}^{2}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}\right),$ (1)
The matrix of estimated channels at the AP $m$ is denoted by $\mathbf{\hat{G}}_{m}=[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]\in\mathbb{C}^{A\times T}$. Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent. As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linear dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the full rank matrix $\mathbf{\bar{G}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is designed as:
$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},$
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]\in\mathbb{C}^{L_{p}\times L_{p}}$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:

### v2

@@ Abstract
A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures. This paper introduces a decentralized decoding architecture that fundamentally breaks from the LSFD, by allowing each access point (AP) to calculate interference-suppressing local weights independently and apply them to its data estimates before transmission. Furthermore, two generalized local zero-forcing (ZF) frameworks, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each AP adaptively and independently determines its combining strategy through a local sum spectral efficiency (SE) optimization that classifies user equipments (UEs) as strong or weak, eliminating the fixed thresholds used in the PFZF and PWPFZF schemes. To enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot. The closed-form SE expressions corresponding to the proposed schemes are derived. Numerical results show that the proposed schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overheads and computation costs with lower performance penalty compared to them.
@@ I Introduction
Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee uniform service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [3, 4]. A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
@@
Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs. Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments. To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE. Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels. These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios. These centralized and partial centralized combining schemes follow the fully centralized architecture as shown in Fig. 1, which has very high computational cost and fronthaul overhead, as for all UEs, the channel needs to be estimated at the CPU with the cooperation of all or a subset of APs.
@@
Several distributed combining approaches have been proposed in the literature. Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference. To improve upon the MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance. However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived. As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult. In [1, 3], the authors follow a simple decoding architecture as in Fig. 1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality with respect to centralized decoding.
@@
The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form SE expressions. In these schemes, each AP classifies the UEs into strong and weak groups, generates one combining vector per orthogonal pilot: the ZF for strong UEs and the MR for weak UEs, and assigns scaled versions to individual UEs based on their estimated channel variances [13]. The fundamental limitations of the PFZF and PWPFZF schemes stem from two factors. First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources. Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
@@
The large-scale fading decoding (LSFD) as shown in Fig. 1, is employed at the CPU to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14]. In the optimal LSFD approach [13], each AP must transmit both channel and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16]. To alleviate this burden, partial LSFD is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process. However, partial LSFD still requires matrix inversions of the same dimension as optimal LSFD and transmission of the channel estimates to the CPU. As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of the LSFD-based architectures in ultra-dense D-mMIMO deployments.
(a) FullyCentralized(b) LSFD(c) SimpleDecoding(d) Proposed DecentralizedDecoding APs CPU All Data Global Combining Global Estimates APs Local Combining CPU Local estimates $\hat{x}_{mt}$Channel estimates LSFD Weights Final Estimates APs Local Combining CPU $\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights $a_{mt}$ CPU $a_{mt}\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}a_{mt}\hat{x}_{mt}$
Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Proposed decentralized decoding.
@@ I-A Contributions
This work addresses some of the above mentioned issues. Our principal contributions are as follows:
@@
Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig. 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead. Central to this architecture is the design of interference-suppressing local weights, computed independently by each AP using only its local channel information. These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator. This eliminates the need for sharing any channel information with the CPU. This transition from a coordinated to a truly decentralized architecture is the key enabler for the practical, large-scale deployment.
@@
Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Unlike the conventional PFZF and PWPFZF schemes, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy. Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR, thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
@@
Novel Decentralized Local Optimization for Pilot Partitioning: We also propose novel decentralized optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum SE metric. This pilot-level optimization is solved using projected gradient ascent optimization, maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
@@
Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable SE of the proposed system. This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
@@
Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good SE with reduced computation cost with the state-of-art baselines. The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul overhead and computational costs in comparison to the LSFD framework and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
@@
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions. Section IV provides a rigorous performance evaluation. Finally, Section V concludes the paper and outlines potential directions for future work.
@@
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
@@ II System Model
The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources. Every AP is linked to the CPU by a fronthaul link, so that the network can operate together and process signals. The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$. We suppose that $\textbf{h}_{mt}\sim\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
@@ II-A Channel Estimation
A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of UE $t$. At the $m$-th AP, the received signal is:
$\textbf{y}^{pilot}_{m}=\sum\nolimits_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission. The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
where $\theta_{mi_{t}}{=}\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$. The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}{=}{\mathbf{g}}_{mt}{-}\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,(\beta_{mt}{-}\gamma_{mt})\textbf{I}_{A})$, where
$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=p^{p}_{t}L_{p}\beta_{mt}^{2}\theta^{-1}_{mi_{t}},$ (1)
Remark 1: The channel estimates of UEs sharing the same pilot sequence are linearly dependent. As a result, an AP cannot fully distinguish among these UEs, leading to residual pilot contamination.
The matrix of estimated channels at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$. Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent. As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the matrix $\mathbf{\bar{G}}_{m}$ is designed as:
$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},\ \ \in\mathbb{C}^{A\times L_{p}}$
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as: